# MigGPT: Harnessing Large Language Models for Automated Migration of Out-of-Tree Linux Kernel Patches Across Versions

### v1

@@ 6.2 Performance Evaluation (RQ1)
As shown in Table 2, in the first step of the two-step strategy–retrieving target code snippets from new versions of the Linux kernel–MigGPT exhibits a significant advantage in the subtask of retrieving target code. Specifically, when paired with a high-performance LLM like GPT-4-turbo, MigGPT achieves a language matching precision of 96.25% for Type 1 samples, significantly outperforming standalone GPT-4-turbo (65.00%). Overall, MigGPT attains an average semantic matching precision of 83.89% across all sample types, marking a 23.15% relative improvement.
@@
As shown in Table 3, in the second step of the two-step strategy—-generating code snippets for new versions of out-of-tree kernel patches—-MigGPT outperforms vanilla LLMs, achieving a 49.26% higher average migration semantic matching precision, a 220% relative improvement. Performance gains increase with the underlying LLM’s capabilities, reaching 80.00% with MigGPT-augmented GPT-4-turbo and 84.44% with MigGPT-augmented DeepSeek-V3.
@@
MigGPT excels in CVE patch backporting task.
We also evaluate the performance of MigGPT in the context of CVE patch backporting. As illustrated in Table 4, MigGPT demonstrates superior performance compared to existing patch backporting methods.
@@
Method Patch SyDIT FixMorph GPT-4-turbo DeepSeek-V3 vanilla MigGPT vanilla MigGPT Accuracy 36% 28% 75% 85% 91% 87% 92%
@@ 6.3 Ablation Study (RQ2)
We conduct an ablation study to evaluate the impact of the four units in MigGPT: CFP, Retrieval Augmentation Module, Retrieval Alignment Module, and Migration Augmentation Module (details in App. E.3). Figure 6 presents the outcomes of four tested variants on our benchmarks. Among these, MigGPT consistently outperforms the ablation baselines in both retrieval and migration tasks. Meanwhile, we perform an ablation study on hyper-parameter $m$ in Algorithm 1 with MigGPT-GPT-4-turbo, which controls the total query time of the Retrieval Augmentation Module. As shown in Figure 6, $m=3$ is suitable for both Type 1 and Type 2 examples.
@@ 6.4 Failure Analysis (RQ3)
We evaluate MigGPT’s robustness by analyzing failed migration cases (not human-matched) across various samples, measuring line edit distances (insertions, deletions, modifications) between MigGPT’s incorrect outputs and human-corrected results (see App. E.1). As shown in Table 5, 41% of MigGPT’s errors require fewer than three lines of modification to align with correct results, demonstrating its potential to aid in out-of-tree kernel patch migration.
Table 5: The number of line edit distances between failed cases and human-migrated outcomes in migration task. “$3\leq\text{dis}<6$” denotes a line edit distance of at least 3 but less than 6.
@@ 7 Conclusion
This study explores the migration of out-of-tree kernel patches in the Linux kernel across various versions. Our proposed benchmark reveals that LLMs struggle with incomplete code context understanding and inaccurate migration point identification. To address these issues, we propose MigGPT, an automated tool for migrating Linux kernel downstream modules. Our evaluation highlights MigGPT’s effectiveness and potential to advance this field.
Impact Statement
@@ A.1 Collection
The migration examples in our benchmark are derived from three open-source out-of-tree kernel patch projects: RT-PREEMPT (Kernel, 2013), HAOC (HAOC, 2024) and Raspberry Pi kernel (Pi, 2018). Notably, RT-PREEMPT’s latest version has been integrated into the mainline Linux kernel for maintenance and no longer exists as an out-of-tree kernel patch. However, this does not impede our utilization of it for research on automated migration and maintenance of out-of-tree kernel patches.
@@ D.2 Examples of Retrieval Augmentation Module
For challenge 1, we construct a “Namespace Prompt” to specify the signatures of the code snippet $s_{\text{old}}$. By constructing the code fingerprint structure $\text{CFP}_{\text{old}}$ from $s_{\text{old}}$ as shown in Figure 8, we can extract FuncDef statements that contain the code signatures (Figure 13), thereby generating a “Namespace Prompt” that describes these signatures. Consequently, the LLM will focus its attention on the function definition __pte_free_tlb rather than on the similar function definition __pmd_free_tlb. This Namespace prompt enhances the LLM’s ability by providing a precise description of the target code, allowing the LLM to focus more accurately on the relevant code snippet and improving the precision of the retrieval.
@@ D.3 Examples of Migration Augmentation Module
The migration augmentation module is primarily designed to address challenge 4. Specifically, as shown in Figure 11, we conduct a comparative analysis between the code fingerprint structures $\text{CFP}_{\text{old}}$ and $\text{CFP}^{\prime}_{\text{old}}$ of the code snippets to ascertain that there are two primary migration points. The first point is located after the comment statement Tial call offset... and before the macro definition statement #define PROLOGUE_OFFSET.... The second point is situated after the statement const struct bpf_prog... and before the statement const int idx0=ctx->idx. By constructing the “Location Prompt”, we enable the LLM to precisely locate the migration points, thereby successfully completing the task of migrating and maintaining the out-of-tree kernel patch.
D.4 Prompts
@@
Appendix E Settings
E.1 Line Edit Distance
@@
$D(A,B)=\begin{cases}\max(n,m)&\text{if }\min(n,m)=0,\\ \min\begin{cases}D(\text{prefix}(A,n-1),B)+1,&\text{(deletion)}\\ D(A,\text{prefix}(B,m-1))+1,&\text{(insertion)}\\ D(\text{prefix}(A,n-1),\text{prefix}(B,m-1))+\mathbb{I}(a_{n}\neq b_{m})&\text{(substitution)}\end{cases}&\text{otherwise.}\end{cases}$
@@
E.2 Threshold of CodeBLEU
CodeBLEU (Ren et al., 2020) is an automated metric designed to evaluate the quality of code generation, specifically tailored for tasks involving the generation of programming code. By integrating both syntactic and semantic features of code, CodeBLEU provides a similarity score ($[0,1]$) between two code snippets. We employ CodeBLEU as a measure of “semantic match” and investigate the alignment between CodeBLEU-based “semantic matches” and “human matches” across various thresholds. As illustrated in Figure 15 and Table 7, we identify a threshold of 0.9 as optimal for our proposed benchmark, ensuring a high degree of consistency between “semantic matches” derived from CodeBLEU and those determined by human evaluation.
@@ E.3 Variant of MigGPT
    MigGPT-No-Retrieval-Augmentation: the Retrieval Augmentation Module is deactivated, causing no constraint on the namespace of code snippets.

### v2

@@ 6.2 Performance Evaluation (RQ1)
As shown in Table 2, MigGPT exhibits a significant advantage in the subtask of retrieving target code. Specifically, when paired with a high-performance LLM like GPT-4-turbo, MigGPT achieves a human matching precision of 96.25% for Type 1 samples, significantly outperforming standalone GPT-4-turbo (65.00%). Overall, MigGPT attains an average semantic matching precision of 84.55% across all sample types, marking a 20.53% relative improvement.
@@
As shown in Table 3, MigGPT outperforms vanilla LLMs, achieving a 73.12% higher average migration semantic matching precision, a 45.61% relative improvement. Notably, due to Llama-3.1-8B’s weak base performance, MigGPT-augmented Llama-3.1-8B underperforms–yet still achieves 44.44% higher semantic matching precision than vanilla Llama-3.1-8B. Additional results for other LLMs are provided in App. E.
@@
Table 4: The compilation success rate MigGPT compared to vanilla LLMs.
Method GPT-4-turbo DeepSeek-V3 Vanilla MigGPT Vanilla MigGPT Rate 16.30% 66.67% (+50.37%) 31.85% 79.26% (+47.41%)
MigGPT performs well on compilation-level metric.
We conducted additional compilation success rate experiments on migrated patches. As shown in Table 4, MigGPT achieves a 48.49% average improvement in compilation success rate over vanilla LLMs.
@@
Method FixMorph TSBPORT PPatHF GPT-4-turbo DeepSeek-V3 vanilla MigGPT vanilla MigGPT Accuracy 24.63% 87.59% 75.12% 85.43% 91.78% 87.12% 92.59%
MigGPT outperformed previous migration efforts.
We also evaluate the performance of MigGPT in the context of CVE patch backporting. Notably, our out-of-tree patch migration task is different from FixMorph, TSBPORT, and PPatHF, which only solve the migration problem, while we target both target code retrieval and migration. As illustrated in Table 5, MigGPT demonstrates superior performance compared to previous patch backporting methods. We also discuss the generalization of MigGPT in App. G.3.
Table 6: The time cost of MigGPT compared to human experts.
Method Expert A Expert B Expert C MigGPT GPT-4-turbo DeepSeek-V3 Time (days) 14.15 10.89 12.51 0.25 0.27
MigGPT is more time-efficient.
We compared MigGPT with three human experts on our out-of-tree patch migration benchmark. As shown in Table 6, MigGPT required only 2.08% of the average time taken by human experts, demonstrating its superior time efficiency.
@@ 6.3 Ablation Study (RQ2)
We conduct an ablation study to evaluate the impact of the four units in MigGPT: CFP, Retrieval Augmentation Module, Retrieval Alignment Module, and Migration Augmentation Module (details in App. F.6). Figure 6 presents the outcomes of four tested variants on our benchmarks. Among these, MigGPT consistently outperforms the ablation baselines. Meanwhile, we perform an ablation study on hyperparameter $m$ in Algorithm 1 with MigGPT-GPT-4-turbo, which controls the total query time of the Retrieval Augmentation Module. As shown in Figure 6, $m=3$ is suitable for both Type 1 and Type 2 examples.
@@ 6.4 Failure Analysis (RQ3)
Table 7: The number of line edit distances between failed cases and human-migrated outcomes in the migration task. $3\leq\text{dis}<6$” denotes a line edit distance of at least 3 but less than 6.
@@
We evaluate MigGPT’s robustness by analyzing failed migration cases (not human-matched) across various samples, measuring line edit distances (insertions, deletions, modifications) between MigGPT’s incorrect outputs and human-corrected results (see App. F.4). As shown in Table 7, 41% of MigGPT’s errors require fewer than three lines of modification to align with correct results, demonstrating its potential to aid in out-of-tree kernel patch migration.
@@ 7 Conclusion
This study explores the migration of out-of-tree kernel patches in the Linux kernel across versions. Our proposed benchmark reveals that LLMs struggle with incomplete code context understanding and inaccurate migration point identification. To address these issues, we propose MigGPT, an automated tool for migrating Linux kernel downstream patches. Our evaluation highlights MigGPT’s effectiveness and potential to advance this field.
@@ A.1 Collection
The migration examples in our benchmark are derived from three open-source out-of-tree kernel patch projects: RT-PREEMPT [19], HAOC [15] and Raspberry Pi kernel [36]. RT-PREEMPT  [^3] https://mirrors.edge.kernel.org/pub/linux/kernel/projects/rt enhances the Linux kernel’s real-time performance for timing-sensitive applications like industrial control and robotics, while Raspberry Pi Linux  [^4] https://github.com/raspberrypi/linux offers a lightweight kernel optimized for embedded systems. HAOC  [^5] https://gitee.com/src-openeuler/kernel/blob/master/ improves kernel security through a "dual-kernel" design, enhancing code behavior, data access, and permission management. These projects are widely adopted in industry and open-source communities, ensuring coverage of critical out-of-tree patch scenarios. Notably, RT-PREEMPT’s latest version has been integrated into the mainline Linux kernel for maintenance and no longer exists as an out-of-tree kernel patch. However, this does not impede our utilization of it for research on automated migration and maintenance of out-of-tree kernel patches.
@@ D.2 Examples of Retrieval Augmentation Module
For challenge 1, we construct a “Structure Prompt" to specify the signatures of the code snippet $s_{\text{old}}$. By constructing the code fingerprint structure $\text{CFP}_{\text{old}}$ from $s_{\text{old}}$ as shown in Figure 8, we can extract FuncDef statements that contain the code signatures (Figure 13), thereby generating a “Structure Prompt" that describes these signatures. Consequently, the LLM will focus its attention on the function definition __pte_free_tlb rather than on the similar function definition __pmd_free_tlb. This Structure prompt enhances the LLM’s ability by providing a precise description of the target code, allowing the LLM to focus more accurately on the relevant code snippet and improving the precision of the retrieval.
@@ D.3 Examples of Migration Augmentation Module
The migration augmentation module is primarily designed to address challenge 4. Specifically, as shown in Figure 11, we conduct a comparative analysis between the code fingerprint structures $\text{CFP}_{\text{old}}$ and $\text{CFP}^{\prime}_{\text{old}}$ of the code snippets to ascertain that there are two primary migration points. The first point is located after the comment statement Tial call offset... and before the macro definition statement #define PROLOGUE_OFFSET.... The second point is situated after the statement const struct bpf_prog... and before the statement const int idx0=ctx->idx. By constructing the “Location Prompt", we enable the LLM to precisely locate the migration points, thereby successfully completing the task of migrating and maintaining the out-of-tree kernel patch.
Algorithm 1 Retrieval of the target code snippet $s_{\text{new}}$
2:Output: $s_{\text{new}}$
3:Generating $\text{CFP}_{\text{old}}$ form $s_{\text{old}}$
4:Preparing RetrievalTaskPromptRetrievalTaskPrompt
5:Preparing ExpertPersonaPromptExpertPersonaPrompt
@@
[…]
11:+RetrievalTaskPrompt+StructurePrompt+RetrievalTaskPrompt+StructurePrompt
12:+AlignmentPrompt+ExpertPersonaPrompt+AlignmentPrompt+ExpertPersonaPrompt
13:while $q<=m$ do
15:  Generating $\text{CFP}_{\text{tmp}}$ from $s_{tmp}$
16:  if $\operatorname{find}(\mathcal{S},s_{tmp})$ then
17:   break
18:  end if
19:  $q\leftarrow q+1$
20:end while
[…]
24:return $s_{\text{new}}$
Algorithm 2 Migration of code snippet $s^{\prime}_{\text{new}}$
2:Output: $s^{\prime}_{\text{new}}$
4:Preparing MigrationTaskPromptMigrationTaskPrompt
5:Preparing ExpertPersonaPromptExpertPersonaPrompt
[…]
9:+LocationPrompt+ExpertPersonaPrompt+LocationPrompt+ExpertPersonaPrompt
[…]
11:return $s^{\prime}_{\text{new}}$
Algorithm 3 The generation of CFP
1:Input: code snippet $s$
2:Output: $\text{CFP}_{s}$
3:Initialize list $\text{CFP}_{s}$
4:for lineline in $s$ do
5:  $Node\leftarrow\operatorname{IdentifyType}(line)$
6:  $(Node.pos,Node.end)\leftarrow\operatorname{IdentifyScope}(line)$
7:  if $Node\in\{\text{FuncDef}\}$ then
8:   $Node.parameter\leftarrow\text{CFP}\left(\operatorname{InternalStatement}(Node)\right)$
9:  end if
10:  if $Node\in\{\text{if},\text{else},\text{while},\text{for},\text{do},\text{switch}\}$ then
11:   $Node.inline\_fuccall\leftarrow\operatorname{Funcall}\left(\operatorname{InternalStatement}(Node)\right)$
12:  end if
13:  $Node.content\leftarrow\operatorname{Content}(Node)$
14:end for
[…]
16:return $\text{CFP}_{s}$
D.4 Algorithm and Prompts
Algorithm 1 and Algorithm 2 respectively illustrate the specific details of target code snippet retrieval and code migration processes. We also present all the prompts utilized by MigGPT. As shown in Figure 14, when retrieving the target code snippet $s_{\text{new}}$, we construct the RetrievalPromptRetrieval\ Prompt to query LLMs. Specifically, we employ TaskPrompt 1Task\ Prompt\ 1 to describe the task and ExpertPersonaPromptExpert\ Persona\ Prompt to standardize the output format of LLMs. Additionally, $StructurePrompt$ and $AlignmentPrompt$ are used to enhance the retrieval capabilities of the LLMs. When generating the migrated code snippet $s^{\prime}_{\text{new}}$, we construct the MigrationPromptMigration\ Prompt to query LLMs. Specifically, we utilize TaskPrompt 2Task\ Prompt\ 2 to describe the task and ExpertPersonaPromptExpert\ Persona\ Prompt to standardize the output format of the large language model. Additionally, $LocationPrompt$ is employed to enhance the migration capabilities of the LLM.
Algorithm 3 illustrates the specific details about generating CFP. The code snippet $s$ is tokenized, and nested scopes (e.g., {,}, #ifdef/#endif) are identified via bracket matching. Critical symbols (e.g., {, }, #ifdef, func()) demarcate code blocks. Function calls (e.g., foo()) are detected through pattern matching on token sequences (e.g., identifier followed by (). Associated functions are identified by analyzing call statements within code blocks, avoiding call graph construction.
@@ D.4 Algorithm and Prompts
Appendix E More Results of MigGPT
We have also tested the performance of MigGPT on DeepSeek-V2.5 [6], as shown in Table 9 and Table 10, MigGPT outperforms the vanilla LLM.
LLM Method Type 1 (80) Type 2 (55) All (135) Average Best Match Semantic Match Human Match Best Match Semantic Match Human Match Best Match Semantic Match Human Match Query Times DeepSeek-V2.5 Vanilla 61.25% 66.25% 62.50% 65.45% 69.09% 67.27% 62.96% 67.40% 64.44% 1.00 MigGPT 95.00% 97.5% 96.25% 87.27% 90.90% 90.90% 91.85% 94.81% 94.07% 1.22
LLM Method Type 1 (80) Type 2 (55) All (135) Best Match Semantic Match Human Match Best Match Semantic Match Human Match Best Match Semantic Match Human Match DeepSeek-V2.5 Vanilla 11.25% 16.25% 18.75% 9.09% 27.27% 21.82% 10.37% 20.74% 20.00% MigGPT 67.50% 80.00% 80.00% 56.36% 74.55% 70.91% 62.96% 77.78% 76.30%
Appendix F Settings
F.1 Platform
Our experiments were conducted on a system equipped with an AMD Ryzen 9 7900X 12-core CPU, 32GB of RAM, running on Ubuntu 22.04.3 LTS.
F.2 The Reliability of Human Match
We strictly adhere to defined steps and principles for Human Match testing, ensuring reliable post-migration code functionality verification:
  • Cross-validation: 5 experienced Linux kernel engineers independently validate results.
  • Triple-blind voting: Each example is evaluated by 3 randomly assigned engineers.
  • Criteria: (1) syntactic consistency, which ensures the preservation of original patch logic through structural adaptations (e.g., variable renaming while maintaining control flow); and (2) semantic correctness, verifying functional equivalence between migrated code and human-generated ground truth patches.
Adherence to these provisions enhances the reliability and credibility of our Human Match metric, ensuring rigorous alignment with established evaluation standards.
F.3 Compilation Success Rate
We conducted additional compilation success rate experiments on migrated patches. For each migration sample, we replace the generated code with MigGPT into the patched version of the new kernel in a containerized environment and attempt to compile the modified kernel. The compilation success rate is the ratio of successfully compiled samples to all types of migration examples.
F.4 Line Edit Distance
@@
$D(A,B)=\begin{cases}\max(n,m)&\text{if }\min(n,m)=0,\\ \min\begin{cases}D(\text{prefix}(A,n-1),B)+1,\\ D(A,\text{prefix}(B,m-1))+1,\\ D(\text{prefix}(A,n-1),\text{prefix}(B,m-1))+\mathbb{I}(a_{n}\neq b_{m})\end{cases}&\text{otherwise.}\end{cases}$ (1)
@@
F.5 Threshold of CodeBLEU
CodeBLEU [37] is an automated metric designed to evaluate the quality of code generation, specifically tailored for tasks involving the generation of programming code. By integrating both syntactic and semantic features of code, CodeBLEU provides a similarity score ($[0,1]$) between two code snippets. We employ CodeBLEU as a measure of “semantic match” and investigate the alignment between CodeBLEU-based “semantic matches” and “human matches” across various thresholds. As illustrated in Figure 15 and Table 11, we identify a threshold of 0.9 as optimal for our proposed benchmark, ensuring a high degree of consistency between “semantic matches” derived from CodeBLEU and those determined by human evaluation.
@@ F.6 Variant of MigGPT
    MigGPT-No-Retrieval-Augmentation: The Retrieval Augmentation Module is deactivated, causing no constraint on the structure of code snippets.
@@
Appendix G Discussion
G.1 Type 1 Migration Sample
Traditional methods like FixMorph [40], SyDIT [28], TSBPORT [49], and PPatHF [34] target single-step migration and cannot solve the target code retrieval problem, while MigGPT tackles both problems, which is harder and more practical. Besides, traditional methods struggle with Type 1 cases due to their reliance on static alignment and predefined transformation rules. As an example shown in Figure 16, when backporting a patch that modifies compute_stats() in the old kernel, FixMorph relies on static alignment (e.g., matching function names like process_data). However, since compute_stats() is now a standalone function in the new kernel, traditional methods like FixMorph cannot generate transformation rules for the split code structure, as its AST differencing assumes code blocks stay within the same function. This illustrates how traditional methods struggle with Type 1’s non-conflicting but structurally divergent changes.
Figure 16: (a) A code snippet of old kernel. (b) The code snippet is refactored into modular functions in the new kernel.
G.2 CFP and Intermediate Representations
A pertinent research question emerges: Could the structural semantics captured through code snippet intermediate representations (IR) provide enhanced contextual signals for guiding LLM-based code migration processes? The answer is negative. While IR-based approaches excel at syntax normalization, our CFP method prioritizes two critical requirements for kernel patch migration:
  • Context Preservation: Kernel patches often contain conditional compilation macros, inline assembly, and annotations stripped during IR generation (e.g., preprocessing eliminates macros). CFP retains these alongside code structure (function signatures, control-flow anchors) to provide LLMs with full migration context.
  • Non-Compilable Code Support: Experimental/unmerged patches (e.g., ARM64-specific optimizations) may fail compilation, making IR extraction impossible. CFP operates directly on source fragments, even for "broken" code in development.
This demonstrates the necessity of employing CFP in MigGPT.
G.3 Generalization of MigGPT
MigGPT demonstrates innovative advancements through its dual strengths of generalizability and domain-specific optimization. While initially designed for out-of-tree kernel patches, its core methodology addresses universal challenges in LLM-based code migration, such as resolving structure conflicts, precisely aligning code boundaries, reconstructing missing context, and establishing migration localization mechanisms, forming a framework extensible to multiple domains (CVE patch backporting [40, 49], forked code porting [34, 51]). The architecture incorporates its Code Fingerprint structure, which embeds Linux kernel-specific optimizations through preserving inline assembly instructions and kernel macro patterns, interpreting kernel-specific comment conventions, and adapting to coding norms like function chaining. This technical design achieves a balance between cross-domain adaptability and deep specialization, with the CFP serving as a modular component that enhances Linux kernel patch migration while maintaining the system’s capacity for expansion into other technical ecosystems. The solution thus enables bidirectional scalability, supporting both horizontal domain transfer and vertical technical refinement.
G.4 Impact Statement